OpenStudy (anonymous):

what is the factored form 6x2 + 5x + 1

OpenStudy (anonymous):



OpenStudy (anonymous):

(2x)((3x+1)

OpenStudy (anonymous):

So first you multiply the first and third factor (6x1=6) and find the factors of the answer (6) that equal the middle term (5), in this case 2 and 3. Now use (a+b)(c+d). ad=2 and bc=3. (2x+3)(x+1)

OpenStudy (anonymous):

If you ever want to ouble check your answer, just foil it back out. (2x+3)(x+1) 2x^2+3x+2x+2 2x^2+5x+2

OpenStudy (anonymous):

Wait dang hold on

OpenStudy (anonymous):

Both your answers are wrong, the first answer when you multiply it back out is 6x2+5x-1 and the other is 2x2+5x+3 and neither of those match the original problem.

OpenStudy (anonymous):

Yes, I do believe I've stated that. Give me a moment, please.

OpenStudy (anonymous):

The easiest way I have found to do these problems is the AC method I believe is what it's called.

OpenStudy (anonymous):

How is that done?

OpenStudy (anonymous):

You have the three parts, ABC...you multiply A and C and take that number and find the factors, then you find the factors that equal B, so in this case 2+3=5 which matches the b section of the problem...rewrite the problem with the factors of 6 in place of 5x so it reads 6x2+3x+2x+1 and then separate into two sides (6x2+2x)(+3x+1) and simplify and the answer is what you have left.....@RavenWinchester

OpenStudy (anonymous):

@bekahyy Is it possible to check the answer in the back of the book?